Sam's brother is 7 years older than Sam his mother is 4 less than 3 times Sam's age their combined age is 63 how old is Sam,his brother,his mother
Answer by ankor@dixie-net.com(22740)

let s = Sam's age
let b = brother's age
let m = Mom's
:
Write an equation for each statement
Sam's brother is 7 years older than Sam
b = s + 7
his mother is 4 less than 3 times Sam's age
m = 3s - 4
their combined age is 63
b + s + m = 63
replace b with (s+7); replace m with (3s-4)
(s+7) + s + (3s-4) = 63
5s + 3 = 63
5s = 63 - 3
5s = 60
s = 12 yrs is Sam's age
:
I'll let you find b and m, see that they add up to 63